OpenStudy (anonymous):

find four concecutiveodd integers such that the sum of the second and third is 19 greater than the fourth

OpenStudy (p0sitr0n):

x x+1 x+2 x+3 (x+1) + (x+2) +19 =( x+3) solve

OpenStudy (anonymous):

let X be the first odd integer, can you write the equations? dont let positron confuse you

OpenStudy (p0sitr0n):

sorry (x+1) + (x+2) =( x+3) +19

OpenStudy (anonymous):

still not right... they are consecutive ODD integers

OpenStudy (p0sitr0n):

19,20,21,22

OpenStudy (anonymous):

20 and 22 are odd?

OpenStudy (p0sitr0n):

concecutiveodd i didnt see the space between

OpenStudy (anonymous):

he is doing it right, but use X and X+2 and X+4 and X+6 as the numbers

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

ok i worked it out and still cant get it

OpenStudy (anonymous):

what did you come up with?

OpenStudy (anonymous):

2x=19

OpenStudy (anonymous):

what was the equation you started with?

OpenStudy (anonymous):

you got it close, but you had an extra X in there somewhere... or forgot to subtract one of them

OpenStudy (p0sitr0n):

x x+2 x+4 x+6 (x+2) +(x+4) = (x+6) +19 x =19?

OpenStudy (anonymous):

x+(x+2) + (x+4)= (x+6 +19

OpenStudy (anonymous):

you are adding the first second and third there... it says add the second and third otherwise, you have it

OpenStudy (anonymous):

so what is the correct equasion.. sorry im confused

OpenStudy (anonymous):

the second + the third = the fourth + 19 (x+2) + (x+4) = (x+6) + 19 you had it, just had an extra X in there

OpenStudy (anonymous):

ok thank you so much:)